# SIGMA: Single Interpolated Generative Model for Anomalies

### v1

@@ I Introduction
This paper is organized as follows: in Section II.1 we describe the dataset used to demonstrate our method. The architecture of the generative model used for our background template is described in  II.2.2. In II.3.1 we describe the slower interpolation method from [3, 7] using flow-matching. The new interpolation method SIGMA is described in Section II.3.2. The model hyperparameters and the training are described in Section II.4. In Section III, we show the results for anomaly detection, and the background template obtained from different methods. Finally in Section IV we present our conclusions and further discussion. In Appendix A, we discuss the importance of using frequency embedding for SIGMA. In Appendix B we analyze the background templates using the classifier test, developed in [38].
@@ III.1 Samples
In Figure 2, we compare the samples from different models to background and data using histograms [^2] Samples from $v^{\rm data}_{\theta}$ have the same normalization as the data distribution. Whereas the samples from $v^{\rm int}_{\theta}$ (linear), $v^{\rm int}_{\theta}$ (context) and $v^{\rm CR}_{\theta}$ have the same normalization as the background distribution. and pull distributions (with respect to the background) [^3] We assume Poisson uncertainty for the pulls., for $N_{\text{sig}}=3000$. We find that pulls for the samples from $v^{\rm data}_{\theta}$ show the presence of signal, similar to the pulls for the data distribution. For the samples from different interpolation methods, higher pull values show significant deviations from the background distribution. All the interpolation methods can smoothen out the signal, but in general $v^{\rm CR}_{\theta}$ seems to produce the best background template, followed by $v^{\rm int}_{\theta}$ (context), and finally $v^{\rm int}_{\theta}$ (linear) being the worst in quality. This is especially clear in the $\Delta R$ distribution, a feature strongly correlated with $m$. Here we see the best pulls from $v^{\rm CR}_{\theta}$, significantly exaggerated pulls from $v^{\rm int}_{\theta}$ (linear), and moderate pull values from $v^{\rm int}_{\theta}$ (context). This shows that $v^{\rm int}_{\theta}$ (context) better models the correlation in $m$ as compared to $v^{\rm int}_{\theta}$ (linear). In Appendix B, we further analyze the samples using the classifier test.

### v2

@@ I Introduction
This paper is organized as follows: in Section II.1 we describe the dataset used to demonstrate our method. The architecture of the generative model used for our background template is described in  II.2.2. In II.3.1 we describe the slower interpolation method from [3, 7] using flow-matching. The new interpolation method SIGMA is described in Section II.3.2. The model hyperparameters and the training are described in Section II.4. In Section III, we show the results for anomaly detection, and the background template obtained from different methods. Finally in Section IV we present our conclusions and further discussion. In Appendix A, we discuss the importance of using frequency embedding for SIGMA. In Appendix B we analyze the background templates using the classifier test, developed in [38]. In Appendix C, we discuss the background template quality for SIGMA for different Signal Regions.
@@ III.1 Samples
In Figure 2, we compare the samples from different models to background and data using histograms [^2] Samples from $v^{\rm data}_{\theta}$ have the same normalization as the data distribution. Whereas the samples from $v^{\rm int}_{\theta}$ (linear), $v^{\rm int}_{\theta}$ (context) and $v^{\rm CR}_{\theta}$ have the same normalization as the background distribution. and pull distributions (with respect to the background) [^3] We assume Poisson uncertainty for the pulls., for $N_{\text{sig}}=3000$. We find that pulls for the samples from $v^{\rm data}_{\theta}$ show the presence of signal, similar to the pulls for the data distribution. For the samples from different interpolation methods, higher pull values show significant deviations from the background distribution. All the interpolation methods can smoothen out the signal, but in general $v^{\rm CR}_{\theta}$ seems to produce the best background template, followed by $v^{\rm int}_{\theta}$ (context), and finally $v^{\rm int}_{\theta}$ (linear) being the worst in quality. This is especially clear in the $\Delta R$ distribution, a feature strongly correlated with $m$. Here we see the best pulls from $v^{\rm CR}_{\theta}$, significantly exaggerated pulls from $v^{\rm int}_{\theta}$ (linear), and moderate pull values from $v^{\rm int}_{\theta}$ (context). This shows that $v^{\rm int}_{\theta}$ (context) better models the correlation in $m$ as compared to $v^{\rm int}_{\theta}$ (linear). In Appendix B, we further analyze the samples using the classifier test. In Appendix C, we show the results for the sample quality for different SRs.
@@ III.2 Anomaly detection performance
III.3 Timing Comparison
In Table 1, we compare the total time taken for generating background templates for different methods. The background template generation method in [7, 3, 64, 22] uses normalizing flows, which takes 3 hours to train. The flow-matching version introduced in this paper, $v^{\rm CR}_{\theta}$, offers a significant speedup to around to 30 mins per SR. The training time for the base model of CURTAINS4F4, which also uses normalizing flows, is around 3 hours, and the lighter flow for each SR trains in 7 mins [26]. Using optimal transport, RAD-OT offered a faster alternative to CURTAINS4F4 at 10 mins per SR. Finally, our method SIGMA, trains the initial flow-matching model for 30 mins and subsequent sampling takes 30 seconds per SR. Hence SIGMA offers significant gains in computational efficiency compared to other methods.
Method Model Timing
CATHODE/ANODE Normalizing Flows 3 hours per SR
CATHODE/ANODE Flow Matching 30 mins per SR
CURTAINS4F4 Normalizing Flows 3 hours
(base model)
+7 mins per SR
RAD-OT Optimal Transport 10 mins per SR
SIGMA Flow Matching 30 mins
+30 secs per SR
Table 1: Timing comparison of different methods.
@@ Appendix B Classifier Test
Figure 7: The figure shows the classifier AUC for background template quality in Signal Regions centered at $m_{JJ}=\left(3.5-0.1\cdot(5-n)\right)~{}\text{TeV}$, for $n=1,\ldots,10$. Classifiers are trained between true background samples in SR, and background template obtained from $v^{\rm int}_{\theta}$ (context). SIGMA is robust across different SRs in the resonant feature.
Appendix C Sample quality across different SRs
In the main body of the paper we focused on the SR $m_{JJ}\in[3.3,3.7]$ TeV that is centered on the LHCO R&D signal with $m_{JJ}=3.5$ TeV. Here we explore the quality of generated samples interpolating into other SRs that do not contain the signal. Taking the same FM model trained on the data which contains 1k signal events as in the main body of the paper, we interpolated it into overlapping SRs of width $400$ GeV, centered at $m_{JJ}=\left(3.5-0.1\cdot(5-n)\right)~{}\text{TeV}$, for $n=1,\ldots,10$. $n=5$ corresponds to the SR considered in the main body of the paper.
For each SR, we trained classifiers between the true background samples (in SR) vs the background template obtained from $v^{\rm int}_{\theta}$ (context). We used the classifier AUC as a metric to assess the background template quality. As shown in Figure 7, we find that the classifier AUCs for all the SRs are close to $\sim 0.5$. This shows that SIGMA is robust across different SRs.